Given, tan(90°–A) = cot(90°–B). What is the value of sin(A–B) ?

Given,
tan(90°–A) = cot(90°–B).
What is the value of sin(A–B) ?

Grade:10

2 Answers

Arun
4 years ago
Dear student
 
As tan (90 -A) = cot (90 -B)
 
Cot A = tan B
Cos A/ sin A - sinB /cos B = 0
 
Cos (A +B) = 0
 
Hemce A +B = 90°
Vikas TU
4 years ago
Dear student 
The explanation is incorrect , 
Cos A/ sin A = sinB /cos B 
CosA CosB – SinASinB = 0 
So , A+ B = 90 
A = 90-B 
A-B = 90-2B 
Sin(A-B) = Sin (90-2B) 
Sin(A-B) = Cos2B 
hope this helps